OpenStudy (mckenzieandjesus):

. Complete. y2 + 15y + 56 = (y + 7)(y +mc001-1.jpg) –7 8 7 –8

OpenStudy (mckenzieandjesus):

the mc001-1.jpg is blank

OpenStudy (anonymous):

I could've been able to solve it if it didn't had those extra stuffs at the end

OpenStudy (mckenzieandjesus):

I said it is blank thats the missing number

OpenStudy (mckenzieandjesus):

its y+ [blank]

OpenStudy (mckenzieandjesus):

@welshfella

OpenStudy (anonymous):

lol...ok let me see if I can do it

OpenStudy (mckenzieandjesus):

i figured it out -7

OpenStudy (anonymous):

Step by step solution : Step 1 : Raise c to the 1st power Exponentiation : Equation at the end of step 1 : (((y2)+15y)+56)-((y+7)•((y+(m•c))-jpg)) = 0 Step 2 : Raise y to the 2nd power Exponentiation : Equation at the end of step 2 : ((y2 + 15y) + 56) - (y + 7) • (y + mc - jpg) = 0 Step 3 : Simplify y2+15y+56 - (y+7)•(y+mc-jpg) 3.1 Factor -ymc+yjpg+8y-7mc+7jpg+56 Try to factor this 6-term polynomial into (2-term) • (3-term) Begin by splitting the 6-term into two 3-term polynomials: -ymc+yjpg+8y and -7mc+7jpg+56 Next simplify each 3-term polynomial by pulling out like terms: -y • (mc-jpg-8) and -7 • (mc-jpg-8) Note that the two simplified polynomials have mc-jpg-8 in common Now adding the two simplified polynomials we get (-7-y) • (mc-jpg-8) Which is the desired factorization. Equation at the end of step 3 : (-y - 7) • (mc - jpg - 8) = 0 Step 4 : Solve (-y-7)•(mc-jpg-8) = 0 Theory - Roots of a product : 4.1 A product of several terms equals zero. When a product of two or more terms equals zero, then at least one of the terms must be zero. We shall now solve each term = 0 separately In other words, we are going to solve as many equations as there are terms in the product Any solution of term = 0 solves product = 0 as well. Solving a Single Variable Equation : 4.2 Solve : -y-7 = 0 Add 7 to both sides of the equation : -y = 7 Multiply both sides of the equation by (-1) : y = -7 Solving a Single Variable Equation : 4.3 Solve mc-jpg-8 = 0 In this type of equations, having more than one variable (unknown), you have to specify for which variable you want the equation solved. We shall not handle this type of equations at this time. One solution was found : y = -7

OpenStudy (anonymous):

please tell me I didn't just embarrassed myself in a public thing...but I hope i'm right